Can someone help me with parts a, b, and c please
Problerr 1 125) Assume that oach of your call not oblaining a busy signal. Assume that your cas ls to a popular radio station has a probability of 0.02 of connecting, that is, of are independent. a) What is the probability tiat your tenth cell is the first one to connect? b) What is the probability that it requires more than five calls for you to connect? c) What is the mean number of calls needed to connect? Page 2 of 6
Answer #1

a) P(X = z) = (1-p)-1p P(X = 10) = (1-0.02)10-10.02 0.016675 0.017 P(X-10) 0.0167 b) P(X > 5)-1-P(X 5) = 1-Fx(5) Recall that